OpenStudy (anonymous):

2a/7 = 2/3 help me with this please!

OpenStudy (anonymous):

i got 6a= 14 ?

OpenStudy (anonymous):

multiply both sides by 7 & then by 3: 3*2a=7*2 6a=14 a=14/6 a=7/3 are you good?

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

which is a= 2and1/3

OpenStudy (anonymous):

thank you!